OpenStudy (anonymous):

find the exact value of sin^-1 (1/2),-Π/2 ≤ x ≤ Π/2

OpenStudy (anonymous):

\[x=\sin^{-1}\frac{1}{2}\] \[\sin x=\frac{1}{2}\] so x=pi/3 for the given range

OpenStudy (anonymous):

can you give me the answer it's my test ? i have to summit tomaroo

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

sorry it is: x=pi/6 (30degrees)

OpenStudy (anonymous):

what do i put for answer?

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

thank you very much can i ask you one more it's similar ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sketch the graph of y = 3sec (x - Π/2) for -2Π ≤ x ≤ 2Π

OpenStudy (anonymous):

help me please

OpenStudy (anonymous):

what is cos(-x) ?

OpenStudy (anonymous):

1/y i think

OpenStudy (anonymous):

cos(-x)=cosx and so is sec(-x)=secx and what is cos(pi/2 - x)?

OpenStudy (anonymous):

is it 0.9997468666 am i right ?

OpenStudy (anonymous):

cos(pi/2 - x)=sinx

OpenStudy (anonymous):

i think you need to clear your basics first

OpenStudy (anonymous):

ow

OpenStudy (anonymous):

i am totaly dumb on math

OpenStudy (anonymous):

i do try hard but it just don't go through my head

OpenStudy (anonymous):

in which class are you?

OpenStudy (anonymous):

precalculus i don't know y i has the class even

OpenStudy (anonymous):

i mean in which grade are you?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

then you can make it,there is nothing too hard in maths

OpenStudy (anonymous):

just you have to make visibility in concepts

OpenStudy (anonymous):

i guess i suck at that

OpenStudy (anonymous):

are you on college?

OpenStudy (anonymous):

can you give me answer of that question it's the end of quater tomaroo i need to pass this test inorder to pass on class? please